OpenStudy (anonymous):

if tan (a+b)=x,tan(a-b)=y ,then what would be the value of tan2a?

OpenStudy (nikvist):

\[\tan{2a}=\tan{((a+b)+(a-b))}=\frac{\tan{(a+b)}+\tan{(a-b)}}{1-\tan{(a+b)}\tan{(a-b)}}=\frac{x+y}{1-xy}\]